# Adaptive Reduced Multilevel Splitting

### v1

@@ 5.1.2 Example #2: reduced basis approximation of a PDE score
We perform standard AMS computations with a score based on this high-fidelity approximation $\Psi^{\ast}(x)$ and using $N=900$ particles. By averaging over 10 runs, we obtain after an extremely intensive calculation an empirical approximation reference of the sought rare event probability $p^{*}_{l_{\max}}=\pi(\,{\rm l}{\rm l}_{S^{\ast}>l_{\max}})\approx 1.47e-04$.
@@
Fig. 2: Illustration of the thermal block problem (example 2). High-fidelity solution $\Psi^{\ast}(X)$ of the PDE for $4$ different samples $X$ of the reference log-normal distribution.
=$\mathcal{M}=\{\Psi^{\ast}(x)\in\mathcal{V}_{h}:x\in\mathcal{X}\},$
by evaluating only some of its elements. By selecting a well-chosen set of $K$ snapshots $\{\Psi^{\ast}(x_{i})\}_{i=1}^{K}$, the RB framework approximates $\mathcal{M}$ in a well-chosen linear subspace $\mathcal{V}_{K}\subset\mathcal{V}_{h}$ of much lower dimension $K\ll h$. Using this framework yields i) an approximation $\Psi$ of the high-fidelity solution $\Psi^{\ast}$ which is used to compute the approximation $S(x)=\|\Psi(x)\|_{\infty}$ of the score $S^{\star}(x)$, ii) a so-called a posteriori error bound $E(x)$ which satisfies $|S^{\ast}(x)-S(x)|\leq\|\Psi^{\ast}(x)-\Psi(x)\|_{\mathcal{V}_{h}}\leq E(x)$. The high-fidelity solution, the RB approximation and the a posteriori error bounds are computed via the “pymor” Python® toolbox [18] available at https://github.com/pymor/pymor. Figure 2 shows several $\Psi^{\ast}(x)$ solutions for samples drawn according to the reference distribution.
Let us comment on two important features of the RB approximation in regards to our rare event simulation framework. First, we emphasize that our algorithm enriches the RB approximation space using a variant of the popular week greedy algorithm. Indeed, the latter selects at each iteration the snapshot in the solution set $\mathcal{M}$ which possess the worst a posterior error bound for the current RB approximation. This setting is used in our snapshot selection in the first stage of the algorithm, i.e., in the case where the learning function selects the parameter maximizing the error. It should be noted, however, that snapshot selection differs in the second stage of our algorithm, in which the RB is enriched with a sample drawn according to the importance distribution.
Refer to caption Refer to caption Refer to caption
Fig. 3: Illustration of RB approximations of a rare event (example 2). The figure displays on the left the high-fidelity solution $\Psi^{\ast}(X)$ of the PDE for a parameter $X$ sampled according to the distribution $\propto\,{\rm l}{\rm l}_{S^{\star}>l_{\max}}d\pi$ (simulated by our ARMS algorithm). Images on the right are RB approximation errors related to $S^{(1)}$ and $S^{(K)}$, the initial and final reduced score approximations.
Figure 3 illustrates a high-fidelity approximation related to the rare event (i.e., $S^{\star}(x)>l_{\max}$) obtained by our algorithm. The figure also shows that the initial RB approximation is insufficiently accurate to discriminate this rare event (i.e., $S(x)<l_{\max}$), while at the end of the algorithm the RB approximation is almost indistinguishable from the high-fidelity one.
@@ 5.2 Results
We begin by detailing the experimental setup for evaluation. Our numerical simulations were carried out using an implementation of the ARMS algorithm in the Julia language, available at https://gitlab.inria.fr/pheas/arms. ARMS is configured as follows: a mutation proposal based on an Ornstein-Uhlenbeck [14] adaptive state movement, a number of particles $N$ is set by default to $31e$ 3 and $21e$ 2 in the first and second examples respectively and explicitly given otherwise, a quantile $\theta=0.3$ (i.e., a maximum number of kills in the selection steps of $M=0.3*N$), the number of hits of the rare event is $J=5$, the number of Markov transitions at each mutation step is set to $21e$ 2 and $13e$ 1 in the first and second examples respectively.
We compute $R$ independent ARMS runs and plot the median and quantiles related to the IS estimate $\hat{p}_{l_{\max}}^{H,N}$ and to the AMS estimate $\hat{p}_{l_{\max}}^{N}$ as a function of the number of snapshots $k$. ARMS performance is quantified by evaluating the expected cost as a function of the relative root expected square error. More specifically, after $k$ snapshots, the relative root expected square error is approximated over runs by the empirical average
$\frac{\sqrt{\mathbb{E}(\hat{p}_{l_{\max}}^{H,N}-p^{*}_{l_{\max}})^{2}}}{{p^{*}_{l_{\max}}}}\approx\left(\frac{1}{R}\sum_{i=1}^{R}\frac{(\hat{p}_{l_{\max},i}^{H_{k},N}-{p^{*}_{l_{\max}})}^{2}}{{p^{*}_{l_{\max}}}^{2}}\right)^{1/2},$ (5.1)
where $\hat{p}_{l_{\max},i}^{H_{k},N}$ denotes, at the $i$-th run and after $k$ snapshots, the importance sampling estimate $\hat{p}_{l_{\max},i}^{H_{k},N}$ if $H_{k}:=k-K_{\min}>0$ and $\hat{p}_{l_{\max},i}^{H_{k},N}=0$ otherwise. After $k$ snapshots, the expected cost (in units of the cost of evaluating $S^{\ast}$) is approximated over the runs by the empirical average
[…]
with $\textrm{gain}^{(k^{\prime})}$ representing the ratio of the cost of evaluating the reduced score $S^{(k^{\prime})}$ to the cost of the true score $S^{\ast}$, and where $N^{(k^{\prime})}_{i}$ represents the number of evaluations of the reduced score at the $k^{\prime}$-th iteration and $i$-th execution of the ARMS algorithm.
Please note that for the first example, $\textrm{gain}^{(k^{\prime})}$ will be set artificially smaller to emulate more realistic cases in which the cost of evaluating the surrogate is indeed cheaper than that of the true model (see Figure 5).
We will compare the performance of ARMS with the empirical average of the estimates obtained by performing $R$ standard AMS with the true score $S^{\ast}$ with a number of particles $N$ in the range $[1e2,1e4]$.
5.2.1 Example #1
$e=42*e_{\min}$ $e=e_{\min}$
Refer to caption Refer to caption
$e=28*e_{\min}$
Refer to caption Refer to caption
Fig. 4: Influence of the entropic criterion threshold $e$ (example 1). Upper and lower left plots: median (in white) and quantiles (in shades of blue-grey) related to the IS estimate (4.10) as a function of the number of snapshots (equal to the number of iteration $k$ of the algorithm) for $3$ different values of the worst log cost threshold $e$. The red line denotes the true rare event probability. Lower right plot: expected cost (5.2) as a function of the relative root expected square error (5.1) (grey curves). Comparison to the performance of standard AMS with $S^{\ast}$ and $N\in[1e2,1e4]$ (blue curve).
Fig. 5: Performance with respect to the reduced score gain (example 1). Expected cost (5.2) as a function of the relative expected squared error (5.1) for different values of the ratio between the cost of evaluating $S^{(k)}$ and $S^{\ast}$ (grey curves) and for the threshold $e=28e_{\min}$. Comparison to the performance of standard AMS with $S^{\ast}$ and $N\in[1e2,1e4]$ (blue curve).
no bridge, $\epsilon=0$, $N=1e3$ bridge, $\epsilon=0$, $N=1e3$
Refer to caption Refer to caption
bridge, $\epsilon=0$, $N=2e4$
Refer to caption Refer to caption
Fig. 6: Bridging performance (example 1). Upper and lower left plots: median (in white) and quantiles (in shades of blue-grey) related to the IS estimate (4.10) as a function of the number of snapshots for $3$ different setups (update stopping criterion $\epsilon$, bridging simulation or not, number of samples $N$). Entropic criterion set to $e=28*e_{\mathrm{min}}$ and gain between surrogate and exact is assumed to be $25$. The red line denotes the true rare event probability. Lower right plot: expected cost (5.2) as a function of the relative root expected square error (5.1) for the various setups. Comparison to the performance of standard AMS with $S^{\ast}$ and $N\in[1e2,1e4]$ (blue curve) and to the performance of the reduced-based estimate (4.12) (green curve).
We perform $R=100$ runs of the ARMS algorithm, each one of them targeting $p_{l_{\max}}^{\ast}\approx 2.2e-8$ with a budget of $K=300$ snapshot evaluations.
@@ 5.2.1 Example #1
Figure 4 evaluates the influence of the threshold $e$ on the entropic criterion of Section 2.2. The three quantiles show the distribution of the IS estimate as a function of the number of snapshots for different values of the entropy criterion threshold $e$. As expected, we note that too high, a threshold produces over 40% of erroneous IS estimates, associated with zero probability regardless of the number of reduced score updates. Although the remaining 60% of estimates nevertheless become accurate after around 60 updates of the reduced score, this demonstrates that increasing the critical level too quickly tends to concentrate particles around unimportant local minima. This results in sampling snapshots and refining the reduced score exclusively in regions of state space unrelated to the rare event of interest. To avoid this undesirable scenario, it is prudent to set the threshold to its minimum value $e_{\min}$, which has the disadvantage of delaying the convergence of the IS estimate. Indeed, the plot shows convergence of all estimates towards a non-zero value close to $p_{l_{\max}}^{\ast}$, but with a considerable delay, since in this case about a hundred more updates of the reduced score are required. A well-tuned value between these extrema achieves convergence after about 70 snapshots, with a characterization of the distribution of IS estimates comparable to the safe distribution obtained with the minimum value.
@@
This trend is confirmed by plotting the expected cost (5.2) as a function of the relative root expected square error (5.1). In the case of a not-too-high threshold $e$, given a certain precision (i.e., after sampling enough snapshots), we observe a gain in expected cost from around a decade onwards, which increases more and more. Clearly, too high a value for $e$ do not achieve comparable accuracy, regardless of the number of snapshots.
@@
The evaluation above focused on the ARMS algorithm in which the simulation is restarted after each snapshot update and the reduced score update stop criterion introduced in Section 2.5 is never triggered ($\epsilon<0$). Figure 6 extends the analysis by evaluating the influence of activating the stop update criterion ($\epsilon\geq 0$) and of bridging the simulations between two iterations, as presented in Section 4.3. The entropic criterion set to $e=28*e_{\mathrm{min}}$. Reference to exact AMS cost is made with a cost gain between surrogate and exact of $25$.
Plotting the predicted cost versus accuracy in the lower right box of Figure 6 shows the considerable savings brought by the bridging procedure.
Some additional remarks on Figure 6 are required.
First, one can remark that, when using the bridging procedure, and for a given sample size, we cannot expect better accuracy for the IS estimator $p_{l_{\max}}^{H,N}$ than for $\hat{p}^{N}_{l_{\max}}$ as both estimates (4.11) and (4.12) depend on the variance of the normalization constant ${Z_{l^{(k)}}^{N,(k)}}$. However, as the same figure shows, because the bridging procedure offers a significant cost saving, we can afford to gain in precision by increasing the number of samples $N$: the curves in yellow in the lower right plot of Figure 6 do show that we obtain a more accurate IS estimate (lower variance and a smaller apparent bias) for a given computational cost.
@@
On the other hand, we notice that enabling the stop-update criterion, even without bridging, induces a spread around the mean of the distribution of the IS estimate in the upper left plot of Figure 6. This expresses (as will show a close look at the expected cost versus accuracy red curve in the lower right plot) the fact that the IS estimator in fact again achieves the accuracy of the reduced score based estimate $p_{l_{\max}}^{N}$ given in (4.12). This latter fact is a consequence of freezing the estimate ${Z_{l^{(k)}}^{N,(k)}}$ as soon as the stop-update criterion is activated and using the latter until the snapshot budget is exhausted. As a result, the variance of the lastest computed normalization constant will affect our IS estimate. The phenomenon does not appear when the stop-update criteria is off (Figure 4) and we restart the simulation as before and calculate a new normalization constant at each iteration (at each iteration). The stop-update criteria is, in fact, simply used to marginally reduce the cost in the bridging procedure, without changing the outcome (not shown in plot).
5.2.2 Example #2
no bridge, $\epsilon=0$, $N=1e2$ bridge, $\epsilon=0$, $N=1e2$
Refer to caption Refer to caption
bridge, $\epsilon=0$, $N=4e3$
Refer to caption Refer to caption
Fig. 7: Performance evaluation (example 2). Upper and lower left plots: median (in white) and quantiles (in shades of blue-grey) related to the IS estimate (4.10) as a function of the number of snapshots for $3$ different setups (update stopping criterion $\epsilon$, bridging simulation or not, number of samples $N$). The red line denotes the true rare event probability. Lower right plot: expected cost (5.2) as a function of the relative root expected square error (5.1) for the various setups. Comparison to the performance of standard AMS with $S^{\ast}$ and $N\in[1e2,2e3]$ (blue curve) and to the performance of the reduced-based estimate (4.12) (green curve).
We perform $R=100$ runs of the ARMS algorithm, each one of them targeting $p_{l_{\max}}^{\ast}\approx 1.5e-4$ with a budget of $K=200$ snapshot evaluations. We set the worst log cost threshold to $e=35\,e_{\min}\approx 0.5$. Overall, the trends of the algorithm for this RB approximation of a PDE score are similar to those of the previous example. Below, we comment on some of the remarkable results.
We analyze in Figure 7 performances using an initial reduced score built on a $5$-dimensional RB approximation. The curves show that about $10$ to $20$ more snapshots are needed to reach convergence, with a median estimate around $p_{l_{\max}}^{\ast}$.
In the case where simulations are not bridged (but we stop updating the reduced score once the worst-case log cost has reached $\epsilon=0$), the distribution of the IS estimate converges after about $50$ more iterations, with some variance but no bias. In the case of bridging, we observe that the distribution of the IS estimate jumps to a constant distribution, with a similar variance but a slight bias. With regard to the latter remark, we suspect that, although the bridging procedure considerably reduces the cost (compared to restarting the simulation in the first iterations of the algorithm), it nevertheless leads to a larger variance of the normalization constant and thus to an apparent bias in the IS estimate. The first remark is simply explained by the fact that the bridging procedure adapts the level for the next iteration of the algorithm to a level so close to $l_{\max}$ that all snapshots are sampled in the rare event, which induces an IS estimate exactly equal to the estimate of the normalization constant ${Z_{l^{(k)}}^{N,(k)}}$.
However, as before, increasing the number of samples $N$ when using bridging yields a reduction of the variance and of the apparent bias while preserving a low expected cost, as shown in the quantile plots and the curves of expected cost versus accuracy in Figure 7.

### v3

@@ 4.1.2 Example #2: reduced basis approximation of a PDE score
As there is no analytic value for the rare event probabilities sought, we perform standard AMS calculations with a score based on this high-fidelity approximation $\Psi^{\ast}(x)$ and using $N=1e3$ particles. Averaging 10 runs, we obtain, after intensive computation, an empirical Monte Carlo approximation for $p^{\ast}=\pi(\,{\rm l}{\rm l}_{S^{\ast}>l_{\max}})$ which is presented in the table above.
$\quad\quad k=1$
Refer to caption Refer to caption Refer to caption Refer to caption
$\quad\quad k=12$
Refer to caption Refer to caption Refer to caption Refer to caption
$\Psi^{\ast}(X^{k+1})$ $\Psi^{(k)}(X^{k+1})$ error $X^{k+1}$
Figure 2: Illustration of the algorithm adaptive behavior (example #2a). The figure displays the high-fidelity solution of the PDE (left), its reduced basis approximation (middle left) and the related absolute error (middle right), evaluated at the snapshot $X^{k+1}$ (right) sampled at iteration $k$ by our ARMS algorithm.
$\quad\quad k=1$
Refer to caption Refer to caption Refer to caption Refer to caption
$\quad\quad k=60$
Refer to caption Refer to caption Refer to caption Refer to caption
$\quad\quad k=120$
Refer to caption Refer to caption Refer to caption Refer to caption
$\quad\quad k=185$
Refer to caption Refer to caption Refer to caption Refer to caption
$\Psi^{\ast}(X^{k+1})$ $\Psi^{(k)}(X^{k+1})$ error $X^{k+1}$
Figure 3: Illustration of the algorithm adaptive behavior (example #2c). The figure displays the high-fidelity solution of the PDE (left), its reduced basis approximation (middle left) and the related absolute error (middle right), evaluated at the snapshot $X^{k+1}$ (right) sampled at iteration $k$ by our ARMS algorithm.
Figure 2 and 3 show typical evolutions along the ARMS algorithm iterations for example #2a and #2c. We observe that in the first iterations of the algorithm, the RB approximation of the high-fidelity solution of the PDE is too rough to characterize the rare event of interest. While the number of iterations increase, RB gain in accuracy, enabling importance sampling of the rare event.
In the context of example #2a, a typical rare event sampled by ARMS corresponds to a diffusion coefficient one decade lower than other diffusion coefficients. To provide a physical interpretation of this rare event, note that each of the $4$ subdomains touches the zero border condition, so that no coefficient configuration can isolate a subdomain. Consequently, increasing the average temperature can only be achieved by slowing down the heat (contributed by the homogeneous source) lost by the system, which corresponds to configurations with low diffusion coefficients. As these configurations lie in the tail of the reference distribution, such events are rare, and of these events, the least rare corresponds to a decrease in one of the $4$ parameters.
In the context of example #2c, a typical rare event sampled by ARMS corresponds to setting the value of most diffusion coefficients touching the zero boundary condition, one decade lower than the value of other diffusion coefficients. The physical interpretation of this rare event is clear: isolating the domain boundary slows the loss of heat contributed by the source, and therefore induces an increase in the average temperature within the domain. Note that such a configuration is extremely rare (on the order of $141e-$ 14) according to the reference distribution, as $11$ coefficients out of $25$ must be sampled in the tail of the reference distribution.
4.2 ARMS setup & evaluation criteria
We now detail the configuration of the ARMS algorithm. The mutation Markov Chain Monte Carlo transition is a Metropolis accept/reject applied on an Ornstein-Uhlenbeck proposal, with an adaptive variance keeping the average acceptance rate over particles in a reasonable range (see e.g.  [15]).
We use a simple scheme to initialize and update the learning parameter $\tau$: in the early iterations of our algorithm ($k\leq K_{j_{0}}$) we set the learning parameter to $\tau^{(k)}=+\infty$ (which corresponds to the weak greedy algorithm setting in the RB framework); once the sampled snapshots have hit $j_{0}$ times the rare event set, we set this parameter to $\tau^{(k)}=0$ for $k>K_{j_{0}}$ (which corresponds to sampling the next snapshots according to the current proposals).
If not explicitly specified in the text, ARMS parameters is configured in accordance with the table below.
$N$ $M/N$ $K$ $T$ $R$ $j_{0}$ $c$
example #1 [2.0e2, 8.0e4] 0.3 300 1e2 100 5 5.0e-2
example #2a [2.5e2, 4.0e2] 0.1 300 3e1 50 2 3.0
example #2b [1.0e2, 4.0e3] 0.3 200 3e1 100 5 5.0e-2
example #2c [1.0e2, 2.0e2] 0.1 300 3e1 50 2 3.0
In the table above, $N$ is the number of particles, $M$ is the number of particles killed at each selection step, $K$ is the snapshot budget, $T$ is the number of Markov transitions at each mutation step, $R$ is the number of independent ARMS runs for each problem, $j_{0}$ is the number of hits before the estimator is triggered and $c$ is the log worst-cost threshold.
We plot the median and quantiles related to the IS estimate $\hat{p}_{IS}$ and to the AMS estimate $\hat{p}_{AMS}$ as a function of the number of snapshots $k$. ARMS performance is quantified by evaluating the expected cost as a function of the relative root expected square error. More specifically, after $k$ snapshots, the expected relative square error is approximated over runs by the empirical average
$\frac{{\mathbb{E}(\hat{p}-p^{\ast})^{2}}}{{p^{\ast}}^{2}}\approx\frac{1}{R}\sum_{r=1}^{R}\frac{(\hat{p}_{r}-{p^{\ast})}^{2}}{{p^{\ast}}^{2}},$ (4.1)
where $\hat{p}_{r}$ denotes, at the $r$-th run and after $k$ snapshots, either AMS estimate $\hat{p}_{AMS}$ given in (3.12), or the importance sampling estimate $\hat{p}_{IS}$ given in (3.11) (with the convention that $\hat{p}_{IS}=0$ if $H=0$). After $k$ snapshots, the expected cost (in units of the cost of evaluating $S^{\ast}$) is approximated over the runs by the empirical average
[…]
with $\textrm{gain}^{(k^{\prime})}$ representing the ratio of the cost of evaluating the reduced score $S^{(k^{\prime})}$ to the cost of the true score $S^{\ast}$, and where $N^{(k^{\prime})}_{r}$ represents the number of evaluations of the reduced score at the $k^{\prime}$-th iteration and $r$-th execution of the ARMS algorithm.
Please note that for the first example, $\textrm{gain}^{(k^{\prime})}$ will be set artificially to emulate realistic cases in which the cost of evaluating the surrogate is indeed cheaper than that of the true model (see Figure 5).
We will compare the performance of ARMS with the empirical average of the estimates obtained by performing $R$ standard AMS with the true score $S^{\ast}$.
$c=42\times c_{\min}$ $c=c_{\min}$
Refer to caption Refer to caption
$c=28\times c_{\min}$
Refer to caption      Refer to caption
Figure 4: Influence of the entropic criterion threshold $c$ (example #1). Upper and lower left plots: median (in white) and quantiles (in shades of blue-grey) related to the IS estimate (3.11) as a function of the number of snapshots (equal to the number of iteration $k$ of the algorithm) for $N=1e3$, a reduced model gain of ×$\times 25$, and $3$ different values of the worst log cost threshold $c$. The red line denotes the true rare event probability. Lower right plot: expected relative square error (4.1) as a function of the expected cost (4.2) (grey curves) for the IS estimate (3.11). Comparison to the performance of standard AMS with $S^{\ast}$ and $N\in[1e2,1e4]$ (blue curve).
Figure 5: Influence of the reduced score gain (example #1). Expected cost (4.2) as a function of the relative expected squared error (4.1) for the IS estimate (3.11), for different values of the ratio between the cost of evaluating $S^{(k)}$ and $S^{\ast}$ (grey curves) and for $N=1e3$ and the threshold $c=28\times c_{\min}$. Comparison to the performance of standard AMS with $S^{\ast}$ and $N\in[1e2,1e4]$ (blue curve).
Figure 6: Bias of the reduced score based estimate (example #1). Median (in white) and quantiles (in shades of blue-grey) related to the AMS estimate (3.12) as a function of the number of snapshots; the red line denotes the true rare event probability. Number of samples $N=1e3$, entropic criterion set to $c=28\times e_{\mathrm{min}}$, no bridging, never stopping the reduced score update ($\epsilon<0$) and gain between surrogate and exact is assumed to be $25$.
no bridge bridge
@@ 4.2 ARMS setup & evaluation criteria
Figure 7: Influence of bridging (example #1). Median (in white) and quantiles (in shades of blue-grey) related to the IS estimate (3.11) (above) and of the cost (4.2) (above), as a function of the number of snapshots when bridging simulation or not; the red line denotes the true rare event probability. Number of samples $N=1e3$, entropic criterion set to $c=28\times e_{\mathrm{min}}$, stopping updates of the reduced score after convergence only in the case of bridging and gain between surrogate and exact is assumed to be $25$. Note that, although imperceptible in the log-log plot at bottom right, the cost increases strictly with the number of snapshots.
Example #1 Example #2b ($L^{\infty}$)
Refer to caption Refer to caption
Example #2a ($L^{1}$) Example #2c ($L^{1}$)
Refer to caption Refer to caption
Figure 8: Performance of ARMS (example #1 &2). Relative expected square error (4.1) as a function of the expected cost (4.2) of the IS estimate (3.11) at the final iteration (once the snapshots budget $K$ is exhausted), for a varying sample size $N$. Curves can be compared to standard AMS estimation using the score $S^{\ast}$, for a varying sample size $N$.
Figure 9: A bad idea: using the stopping criterion without bridging (example #1). Median (in white) and quantiles (in shades of blue-grey) related to the IS estimate (3.11) as a function of the number of snapshots when stopping the reduced score updates ($\epsilon=0$) without bridging; the red line denotes the true rare event probability. Number of samples $N=1e3$, entropic criterion set to $c=28\times c_{\mathrm{min}}$ and gain between surrogate and exact is assumed to be $25$.
@@ 4.3 Results
We begin by analyzing the results obtained with the toy model of example #1. Figure 4 evaluates the influence of the threshold $c$ on the entropic criterion of Section 2.2. The three quantiles show the distribution of the IS estimate as a function of the number of snapshots for different values of the entropy criterion threshold $c$. As expected, we note (upper left) that a too large threshold produces over 40% of erroneous IS estimates, associated with zero probability regardless of the number of reduced score updates. Although the remaining 60% of estimates nevertheless become accurate after around 60 updates of the reduced score, this demonstrates that increasing the critical level too quickly tends to concentrate particles around unimportant local minima. This results in sampling snapshots and refining the reduced score exclusively in regions of state space unrelated to the rare event of interest. To avoid this undesirable scenario, it is possible to set the threshold to its minimum value $c_{\min}$ (upper right), which has the disadvantage of delaying the convergence of the IS estimate. Indeed, the plot shows convergence of all estimates towards a non-zero value close to $p^{\ast}$, but with a considerable delay, since in this case about a hundred more updates of the reduced score are required. A well-tuned value between these extrema (lower left) achieves convergence after about 70 snapshots, with a characterization of the distribution of IS estimates comparable to the safe distribution obtained with the minimum value.
@@
This trend is confirmed by plotting (lower right) the relative root expected square error (4.1) as a function of the expected cost (4.2). In the case of a not-too-high threshold $c_{\min}$ or $28c_{\min}$, given a certain precision (i.e., after sampling enough snapshots), we observe a gain in expected cost from around a decade onwards, which increases more and more. Clearly, a value of $42c_{\min}$ is too high and do not achieve comparable accuracy, regardless of the number of snapshots.
@@
Figure 6 shows that the AMS estimate (3.12) significantly underestimates the true probability, demonstrating that relying solely on reduced scores $S^{(k)}$ (here spline approximations) of the true score $S^{\ast}$ induces a significant bias. Indeed, in our specific example, the non-differentiability of $S^{\ast}$ at the rare event boundary leads to high-frequency oscillations (which tend to decrease with the number of interpolation points, but never disappear completely) in the spline model. The consequence is that a significant proportion of particles in the rare event set $\{x\in\mathcal{X}:S^{\ast}(x)\geq l_{\max}\}$ are to be excluded from the set $\{x\in\mathcal{X}:S^{(k)}(x)\geq l_{\max}\}$, leading to a negative bias. On the contrary, we observe that the IS estimate (3.11) is not biased, independently of the problem of convergence of $S^{(k)}$ to $S^{\ast}$.
The evaluation above focused on the ARMS algorithm in which the simulation is restarted after each reduced score update and in which the stop criterion introduced in Section 2.5 is never triggered ($\epsilon<0$). Figure 7 extends the analysis by evaluating the influence of bridging the simulations between two iterations, as presented in Section 3.3. The entropic criterion is set to $c=28\times c_{\mathrm{min}}$. As shown in figure 7, the gain in computational cost brought about by bridging is offset by an increase in variance. One can remark that, when using the bridging procedure, and for a given sample size, we cannot expect to increase the accuracy for the IS estimate (3.11) indefinitely, contrary to when the simulation is restarted after each reduced score update. Indeed, the estimate depends crucially on the variance of the normalization constant ${Z_{l^{(k)}}^{N,(k)}}$ which generates a residual error (upper right box of Figure 7). In the case of bridging, this residual error will indeed propagate across the iterations according to (3.9) instead of being re-estimated by restarting a fresh simulation at each reduced score update. Nevertheless, the plot of expected squared error versus predicted cost in the upper left-hand panel of figure 8 shows that, in the case of small sample sizes yielding a relative square error above 20%, some saving of order of 2 is achieved by a bridging procedure compared to restarting the simulation, while both procedures are equivalent for larger sample sizes.
@@
On the other hand, we notice that enabling the stopping criterion without bridging, is a bad idea as it induces a spread around the mean of the distribution of the IS estimate as illustrated in Figure 9. Indeed, similarly to the above explanation, the IS estimate will be affected by the apparent bias of the estimates ${Z_{l^{(k)}}^{N,(k)}}$ in (3.11), all identical for indices $k$ above some random time (determined by the stopping criterion). The phenomenon does not appear when the stopping criterion is off (Figure 7) and we restart the simulation as before and calculate a new normalization constant at each iteration. The stopping criterion is, in fact, simply used to marginally reduce the cost in the bridging procedure, without changing the outcome (not shown in plot).
We then move on to example #2, the PDE-based scores and their RB approximations. Overall, the trends of the algorithm for the RB approximation of a PDE score are very similar to those of the previous example. We thus omit the precise evaluation made for the previous example, and jump directly to the analysis of performances in Figure 8. We use an initial reduced score built on a $5$-dimensional RB approximation. In the case of examples #2a and #2b, we observe that about $10$ to $20$ more snapshots (i.e., ARMS iterations) are needed to reach convergence, with a median estimate around $p^{\ast}$. In the case of example #2c, about $170$ to $200$ additional snapshots are needed to reach a good accuracy of the reduced model and converge towards $p^{\ast}$. Indeed, in the latter situation, the parameter space has a much higher dimensionality (25 dimensions versus 4), and learning the reduced model for parameters in the (initially unknown) region where rare events occur is a difficult task performed progressively over the course of ARMS iterations.
In all examples, Figure 8 shows that the proposed ARMS method enables important computational savings as compared to the use of the true model alone. Saving are consistently around a factor $10$ in all cases. As mentioned in section 4.1.2, in the case of the #$\#2b$ example, ARMS uses error estimates that are not hard bounds for the $L^{\infty}$ score. Nevertheless, it is surprising that ARMS manages to accurately estimate the desired probability and achieve performances comparable to those obtained with the $L^{1}$ score.